# Anomalous dimensions from conformal field theory: generalized $\phi^{2n+1}$ theories

### v2

@@ 4.1 Multiplet recombination
$\mathcal{S}(n)=\sum_{\begin{subarray}{c}r+s+t=2n+1\\ r,s,t\in\mathbb{Z}_{+}\end{subarray}}\Bigg{[}$ $\;\frac{1}{r!s!t!(2n+1-r-s)!(2n+1-s-t)!(2n+1-t-r)!}$
×$\quad\times\frac{\Gamma\left(\frac{s+t-r}{2n-1}\right)\Gamma\left(\frac{t+r-s}{2n-1}\right)\Gamma\left(\frac{r+s-t}{2n-1}\right)}{\Gamma\left(\frac{2r}{2n-1}\right)\Gamma\left(\frac{2s}{2n-1}\right)\Gamma\left(\frac{2t}{2n-1}\right)}\;\Bigg{]}\,.$ (97)
Here $\mathbb{Z}_{+}$ denotes the set of positive integers. If we use this input from the diagrammatic method, then the anomalous dimensions of $\phi$ and $\mathcal{J}$ are completely fixed for $k=1$. For $n=1$, there is only one term in (4.1), which is associated with $r=s=t=1$. We verify that the $n=1$ case of (96) matches the Yang-Lee result (23). For $n=2,3,4$, our results for $\gamma_{\phi}$ are in agreement with those of Gracey:2017okb . [^29] However, our results for $\gamma_{\phi^{2}}$ do not agree with Gracey:2017okb . Some disagreement was pointed out by Codello:2017epp as well, and it was shown that the results of Codello:2017epp are supported by the CFT analysis Codello:2017qek .
@@ 4.2 Analytic bootstrap
In this subsection, we verify the multiplet-recombination results using the analytic bootstrap. The multiplet-recombination results are consistent with the leading crossing constraints for arbitrary $\alpha$. Therefore, as in the multiplet recombination approach, we are not able determine $\alpha$ by the leading CFT analysis. As in subsection 2.2, we consider the four-point function
@@
At order $\alpha^{0}$, this analysis is the similar to that in subsection 2.2. The difference is that the scaling dimension of $\phi_{\text{f}}$ takes a different value $\Delta_{\smash{\phi_{\text{f}}}}=\frac{2k}{2n-1}$. We derive the free OPE coefficient
@@
=\;= $\frac{n(2n)!\,\Gamma(2k)\alpha^{2}}{2^{4k-1}\Gamma(k)^{2}(1-2k)_{k}^{2}\left(k\frac{3-2n}{2n-1}\right)_{k}^{2}}\,\bar{z}^{\frac{2k}{2n-1}}(1-\bar{z})^{k\frac{2n-3}{2n-1}}{}_{2}F_{1}\!\left(k,k;k\frac{2n-3}{2n-1}+1;1-\bar{z}\right)+O(\alpha^{3})\,,$ (101)
where we have used the squared OPE coefficient (160) derived in appendix F. We should be more careful, as there are infinitely many terms contributing to the enhanced singularity. A more elegant approach is to use the Lorentzian inversion formula. In the lightcone limit, this leads to an SL($2,\mathbb{R}$) inversion integral Caron-Huot:2017vep ; Alday:2017zzv
@@
where we consider the leading term in the $z\rightarrow 0$ expansion. The double discontinuity is defined as
@@
The calculation of the dDisc associated with (4.2) amounts to multiplying by the factor $2\sin^{2}(\pi k\,\frac{2n-3}{2n-1})$. Substituting this dDisc into the inversion integral (102) and dividing by the free OPE coefficient (100), we obtain the anomalous dimensions of the bilinear operators on the $m=0$ trajectory. The results agree with those from the multiplet recombination, confirming that the multiplet-recombination results are compatible with OPE associativity. In appendix B, the analytic bootstrap has also been carried out at subleading twist, where the multiplet-recombination results are again consistent with crossing symmetry.
@@ 5 Discussion
In this paper we have computed the anomalous dimensions of $\phi$ and bilinear operators $\mathcal{J}$ in the generalized $\phi^{2n+1}$ theories, using the conformal multiplet recombination. For $n=1$, we have also considered the $\phi^{3}$ theory with global symmetry $S_{N+1}$. Our main results are summarized in table 1. We have verified that they are consistent with crossing symmetry using the analytic bootstrap. To the best of our knowledge, new results in the canonical case $k=1$ include the anomalous dimensions of spinning bilinear operators in the Potts model. In the generalized case with a higher-derivative kinetic term, i.e. $k>1$, all the results for the anomalous dimensions and OPE coefficients should be new.
@@
We have considered the $S_{N+1}$ global symmetry, but have restricted to the $\phi^{3}$ case. It might be interesting to study the $S_{N+1}$-symmetric $\phi^{p}$ theories with higher $p$, as they should provide a variety of universality classes. In particular, the canonical $\phi^{4}$ and $\phi^{5}$ theories with $S_{N+1}$ symmetry have been studied in Rong:2017cow ; Codello:2018nbe ; Codello:2020mnt . These theories correspond to the quartic and quintic Potts models. It would also be interesting to investigate their higher-derivative generalizations.
@@
Since the set of anomalous dimensions is fixed up to an unknown parameter $\alpha$ in the $n>1$ theory, another important question is if we can determine $\alpha$ in the $n>1$ theory by conformal symmetry and consistency requirements. In the canonical case $k=1$, we fix the value of $\alpha$ by the input from Codello:2017epp , which is given in (96). Given the complicated summation in (4.1), it seems unlikely that we can determine $\alpha$ by the leading analysis of the multiplet recombination method. As (4.1) shares some similarities with the subleading corrections in the $\phi^{2n}$ theory Codello:2017epp , we expect that $\alpha$ can be determined by a higher order study, which requires a merging of the multiplet recombination and crossing analyses. As a necessary step, we verified that the multiplet-recombination results are consistent with crossing symmetry for the leading corrections.
@@
As the generalized free theory is non-unitary, we expect that the $\phi^{2n+1}$ deformations are non-unitary as well. The nonperturbative bootstrap study of their integer $d$ versions requires bootstrap methods that do not rely on positivity constraints Gliozzi:2013ysa ; Gliozzi:2014jsa ; El-Showk:2016mxr ; Esterlis:2016psv ; Li:2017agi ; Li:2017ukc ; Li:2021uki ; Kantor:2021kbx ; Kantor:2021jpz ; Afkhami-Jeddi:2021iuw ; Laio:2022ayq ; Li:2023tic . In Guo:2023qtt and the present work, we have investigated the $\phi^{2n}$ and $\phi^{2n+1}$ deformations of the generalized free theory, i.e., $\phi^{m}$ interactions with integer power $m\geq 3$.  [^30] Can the generalized $\phi^{m}$ theories be realized by lattice models? A natural extension is to consider non-integer $m$. Some simple examples are the fractional power $m=\frac{m_{1}}{2},\frac{m_{1}}{3},\frac{m_{1}}{4},\dots$, where $m_{1}$ is a large enough integer. The $\phi^{m}$ deformations with irrational power or even complex power are also interesting from the perspective of analytic continuation, as they can connect the standard integer-power cases in the theory space, but we expect that they are non-unitary even for the canonical case of $k=1$. The $\mathcal{PT}$ symmetric theories are particularly relevant because they can satisfy the physical requirement of a bounded-from-below spectrum in a continuous range of power Bender:1998ke ; Bender:2007nj ; Bender:2023cem .  [^31] It is not clear to us how to capture the $\mathcal{PT}$ symmetric CFTs with even power $m$ using the $\epsilon$ expansion, as one needs to be more careful about the quantization scheme and it is not an easy task to handle multidimensional Stokes sectors, i.e., Lefschetz thimbles. See Felski:2021evi and references therein. Furthermore, it seems more natural to use the combination $i\phi$ as the building block for $\mathcal{PT}$ symmetric theories. Recently, the low dimensional $\mathcal{PT}$ symmetric models have been revisited by bootstrap methods in Li:2022prn ; Khan:2022uyz ; Li:2023nip ; John:2023him ; Li:2023ewe ; Li:2024rod , using the equations of motion in the Lagrangian or Hamiltonian formalism. It would be interesting to extend the low dimensional bootstrap methods to higher dimensions. Along the lines of analytic continuation, one may also bootstrap the non-perturbative conformal field theory by the homotopy continuation method Li:2024ggr , where the $\epsilon$ expansion can provide regular boundary conditions near but away from the potentially dangerous Gaussian limit.  [^32] Conformal symmetry implies the existence of zero-norm states associated with specific Gaussian scaling dimensions. Usually, the zero-norm states should decouple to avoid inconsistency, such as divergences in the Gaussian limit.
@@
In addition, the $\phi^{4}$ deformation of a canonical free theory with a boundary or monodromy/line defect has been studied using the multiplet recombination Yamaguchi:2016pbj ; Dey:2020jlc ; Nishioka:2022odm ; Nishioka:2022qmj . An interesting future direction is to consider the $\phi^{m}$ deformations and higher-derivative generalizations, as well as other types of defects. Another direction would be to study CFTs on nontrivial manifolds. The case of real projective space has been carried out for the $\phi^{3}$ and $\phi^{4}$ theory Hasegawa:2016piv ; Hasegawa:2018yqg . It would also be interesting to apply the multiplet recombination method to CFTs at finite temperature.
@@ Appendix A The action of Laplacians
Here $\Delta_{i}$ denotes scaling dimension of $\mathcal{O}_{i}$. The result of the $\Box_{x_{1}}$ action takes the form
@@
=\;= $\left[\frac{h_{123}\left(h_{123}+h_{132}+2-d\right)}{x_{12}^{2}}+\frac{h_{132}\left(h_{123}+h_{132}+2-d\right)}{x_{13}^{2}}-\frac{h_{123}h_{132}\,x_{23}^{2}}{x_{12}^{2}x_{13}^{2}}\right]\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}\,.$ (107)
@@
$C_{i,j}^{a,b}$ $=\left(\frac{a}{2}\right)_{i}\left(\frac{b}{2}\right)_{j}\left(\frac{a+b}{2}-\frac{d-2}{2}+i+j-k\right)_{2k-i-j}\,.$ (110)
@@
×$\times B_{i_{3},j_{3}}C_{i_{3},j_{3}}^{h_{231}+2(j_{2}+k-i_{1}-j_{1}),h_{132}+2(j_{1}+k-i_{2}-j_{2})}x_{12}^{2(i_{3}+j_{3}-i_{1}-i_{2}-k)}$
×$\times x_{13}^{2(i_{2}+j_{2}-j_{1}-j_{3}-k)}x_{23}^{2(i_{1}+j_{1}-j_{2}-i_{3}-k)}\Bigg{]}\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}\,.$ (111)
@@ Appendix B Analytic bootstrap at subleading twist
Let us start with the generalized Yang-Lee edge singularity. We examine the subleading term of the crossing equation (33) in the $z\rightarrow 0$ expansion. On the left-hand side of the crossing equation, both $\mathcal{J}^{(0)}_{\ell}$ and $\mathcal{J}^{(1)}_{\ell}$ contributes to the infinite sum. The lightcone expansion of the conformal blocks are written as
$G_{\mathcal{O}}(z,\bar{z})=z^{(\Delta_{\mathcal{O}}-\ell)/2}k_{h}(\bar{z})+\sum_{i=-1}^{1}c_{i}\,z^{(\Delta_{\mathcal{O}}-\ell)/2+1}k_{h+i}(\bar{z})+O(z^{(\Delta_{\mathcal{O}}-\ell)/2+2})\,,$ (112)
For $\mathcal{O}=\mathcal{J}^{(0)}_{\ell}$, the relevant part is the subleading term in this expansion. Here $c_{i}$ are obtained by requiring that (112) solves the Casimir equation Dolan:2003hv . The results are
@@
$+z^{\Delta_{\smash{\phi}}+1}\left(\sum_{i=-1}^{1}c_{i}k_{h^{(0)}+i}(\bar{z})+\frac{1}{2}\tilde{\gamma}_{\mathcal{J}^{(0)}_{\ell}}\sum_{i=-1}^{1}c_{i}k_{h^{(0)}+i}(\bar{z})\log z+O(\tilde{\gamma}_{\smash{\mathcal{J}^{(0)}_{\ell}}}^{2})\right)$
$+O(z^{(\Delta_{\smash{\mathcal{J}^{(0)}_{\ell}}}-\ell)/2+2})\,,$ (114)
@@
$G_{\mathcal{J}^{(1)}_{\ell}}(z,\bar{z})=z^{\Delta_{\smash{\phi}}+1}\left(k_{h^{(1)}}(\bar{z})+\frac{1}{2}\tilde{\gamma}_{\mathcal{J}^{(1)}_{\ell}}k_{h^{(1)}}(\bar{z})\log z+O(\tilde{\gamma}_{\smash{\mathcal{J}^{(1)}_{\ell}}}^{2})\right)+O(z^{(\Delta_{\smash{\mathcal{J}^{(1)}_{\ell}}}-\ell)/2+1})\,.$ (115)
@@
The enhanced singularity on the right-hand side of the crossing equation is $\Delta_{\smash{\phi_{\text{f}}}}(\frac{1-\bar{z}}{\bar{z}})^{-\Delta_{\smash{\phi_{\text{f}}}}}$, which corresponds to the identity operator contribution. Since $\Delta_{\smash{\phi_{\text{f}}}}=2k$, the asymptotic relation implies that the coefficient in (117) is given by $2S_{-2k}(h^{(1)})$ at order $\epsilon^{0}$. Subtracting the contribution from the leading twist operators $\mathcal{J}^{(0)}_{\ell}$, we find the free squared OPE coefficients
@@
$P_{\phi}\frac{\Gamma(2k)}{\Gamma(k)^{2}}\Bigg{[}\left(\Delta_{\phi}+\frac{\Delta_{\phi}\left(\Delta_{\phi}-2\right)}{4}\right)\sum_{i=0}^{k-1}(-1)^{i+1}\frac{(1-2k)_{i}(k)_{i}}{i!\,(1-k)_{i}}\left(\frac{1-\bar{z}}{\bar{z}}\right)^{i-k}$
$\qquad\qquad\;-\frac{\left(d-2\right)\Delta_{\phi}^{2}}{4\left(d-2\Delta_{\phi}-2\right)}\sum_{i=0}^{k-2}(-1)^{i+1}\frac{(1-2k)_{i}(k+1)_{i}}{i!\,(2-k)_{i}}\left(\frac{1-\bar{z}}{\bar{z}}\right)^{i-k+1}$
@@
where we have only kept the terms of order $\epsilon^{1}$. Using (24), (37) and (39), we find the coefficients in the infinite sum (119). The anomalous dimensions $\tilde{\gamma}_{\smash{\mathcal{J}^{(1)}_{\ell}}}$ are obtained by subtracting the contribution from $\mathcal{J}^{(0)}_{\ell}$. The results agree with those from the multiplet recombination.
@@
which is the $m=1$ version of (78) at leading twist. For the anomalous dimensions, we consider the enhanced singularity given by (B). From this we extract the coefficients in the infinite sum of the form (119). Here we need to account for the $N$-dependent factors in the crossing equations (D)-(D). As a result, the coefficients in the infinite sums have the following relation with those in the Yang-Lee case:
@@
Multiplying by the factor $2\sin^{2}(\pi k\,\frac{2n-3}{2n-1})$, we obtain the corresponding dDsic, as in (4.2). At subleading order in the $z\rightarrow 0$ expansion, we find the anomalous dimensions of $\mathcal{J}^{(1)}_{\ell}$ using the leading-twist anomalous dimension, as well as the free squared OPE coefficients (100) and (124). The results are consistent with those obtained from the multiplet recombination.

### v3

@@ 4.1 Multiplet recombination
$\mathcal{S}(n)=\sum_{\begin{subarray}{c}r+s+t=2n+1\\ r,s,t\in\mathbb{Z}_{+}\end{subarray}}\!\!\;\frac{\Gamma(\frac{s+t-r}{2n-1})\Gamma(\frac{t+r-s}{2n-1})\Gamma(\frac{r+s-t}{2n-1})}{r!s!t!(2n+1-r-s)!(2n+1-s-t)!(2n+1-t-r)!\,\Gamma(\frac{2r}{2n-1})\Gamma(\frac{2s}{2n-1})\Gamma(\frac{2t}{2n-1})}\,.$ (98)
Here $\mathbb{Z}_{+}$ denotes the set of positive integers. If we use this input from the diagrammatic method, then the anomalous dimensions of $\phi$ and $\mathcal{J}$ are completely fixed for $k=1$. For $n=1$, there is only one term in (98), which is associated with $r=s=t=1$. We verify that the $n=1$ case of (97) matches the Yang-Lee result (21). For $n=2,3,4$, our results for $\gamma_{\phi}$ are in agreement with those of Gracey:2017okb . [^37] However, our results for $\gamma_{\phi^{2}}$ do not agree with Gracey:2017okb . Some disagreement was pointed out in Codello:2017epp as well, and it was shown that the results of Codello:2017epp are supported by the CFT analysis Codello:2017qek .
@@ 4.2 Analytic bootstrap
In this subsection, we verify the multiplet-recombination results using the analytic bootstrap. As noted in the generalized Yang-Lee and Potts theories below (44) and (80), the multiplet-recombination results are consistent with the leading crossing constraints for arbitrary $\alpha$. Therefore, as in the multiplet-recombination approach, we are not able determine $\alpha$ by the leading crossing analysis. As in section 2.2, we consider the four-point function
@@
At order $\alpha^{0}$, this analysis is the similar to that in section 2.2. The difference is that the scaling dimension of $\phi_{\text{f}}$ takes a different value $\Delta_{\smash{\phi_{\text{f}}}}=\frac{2k}{2n-1}$. We derive the free OPE coefficient
@@
=\;= $-\frac{n(2n)!\,\Gamma(2k)\alpha^{2}}{2^{4k-1}\Gamma(k)^{2}(k)_{k}^{2}\left(-k\frac{2n-3}{2n-1}\right)_{k}^{2}}\,\bar{z}^{\frac{2k}{2n-1}}(1-\bar{z})^{k\frac{2n-3}{2n-1}}{}_{2}F_{1}\!\left(k,k;k\frac{2n-3}{2n-1}+1;1-\bar{z}\right)+\ldots\,,$ (102)
where we have used the squared OPE coefficient (162) derived in appendix F. We should be more careful, as there are infinitely many terms contributing to the enhanced singularity. Expanding (4.2) in terms of $\frac{1-\bar{z}}{\bar{z}}$, we can derive the large-spin expansion of the anomalous dimensions $\tilde{\gamma}_{\smash{\mathcal{J}^{(0)}_{\ell}}}$ order by order using the asymptotic relation (37). A more elegant approach is to use the Lorentzian inversion formula, which gives the complete result as an analytic function in spin. In the lightcone limit, this leads to an SL($2,\mathbb{R}$) inversion integral Caron-Huot:2017vep ; Alday:2017zzv
@@
where we consider the leading term in the $z\rightarrow 0$ expansion. The double discontinuity (dDisc) is defined as
@@
The calculation of the dDisc associated with (4.2) amounts to multiplying by the factor $2\sin^{2}(\pi k\,\frac{2n-3}{2n-1})$. Substituting this dDisc into the inversion integral (103) and dividing by the free OPE coefficient (101), we obtain the anomalous dimensions of the bilinear operators on the $m=0$ trajectory:
$\tilde{\gamma}_{\mathcal{J}_{\ell}^{(0)}}=-\frac{n^{2}(2n-1)!\,\alpha^{2}}{2^{4k-3}(2k-1)!\left(\ell-k\frac{2n-3}{2n-1}\right)_{2k}}+\ldots\,.$ (106)
The result agrees with (96) at $m=0$ and confirms the compatibility of the multiplet recombination with OPE associativity. In appendix B, the analytic bootstrap is also carried out at subleading twist, where the multiplet-recombination results are again consistent with crossing symmetry.
@@ 5 Discussion
In this paper we computed the anomalous dimensions of $\phi$ and bilinear operators $\mathcal{J}_{\ell}^{(m)}\sim\phi\partial^{\ell}\Box^{m}\phi$ in the generalized $\phi^{2n+1}$ theories, using the conformal multiplet recombination. For $n=1$, we also considered the $\phi^{3}$ theory with global symmetry $S_{N+1}$. Our main results are summarized in table 1. We verified that they are consistent with crossing symmetry using the analytic bootstrap. To the best of our knowledge, new results in the canonical case $k=1$ include the anomalous dimensions of spinning bilinear operators in the Potts model. In the generalized case with a higher-derivative kinetic term, i.e., $k>1$, all of the results for the anomalous dimensions and OPE coefficients should be new.
@@
We considered the $S_{N+1}$ global symmetry, but restricted to the $\phi^{3}$ case. It might be interesting to study the $S_{N+1}$-symmetric $\phi^{p}$ theories with higher $p$, as they should provide a variety of universality classes. In particular, the canonical $\phi^{4}$ and $\phi^{5}$ theories with $S_{N+1}$ symmetry were studied in Rong:2017cow ; Codello:2018nbe ; Codello:2020mnt . These theories correspond to the quartic and quintic Potts models. It would also be interesting to investigate their higher-derivative generalizations.
@@
Since the set of anomalous dimensions is fixed up to an unknown parameter $\alpha$ in the $n>1$ theory, another important question is if we can determine $\alpha$ in the $n>1$ theory by conformal symmetry and consistency requirements. In the canonical case $k=1$, we fix the value of $\alpha$ by the input from Codello:2017epp , which is given in (97). Given the complicated summation in (98), it seems unlikely that we can determine $\alpha$ by the leading analysis of the multiplet recombination method. As (98) shares some similarities with the subleading corrections in the $\phi^{2n}$ theory Codello:2017epp , we expect that $\alpha$ can be determined by a higher-order study, which requires a merging of the multiplet recombination and crossing analyses. As a necessary step, we verified that the multiplet-recombination results are consistent with crossing symmetry for the leading corrections.
@@
As the generalized free theory is nonunitary, we expect that the $\phi^{2n+1}$ deformations are nonunitary as well. The nonperturbative bootstrap study of their integer $d$ versions requires bootstrap methods that do not rely on positivity constraints Gliozzi:2013ysa ; Gliozzi:2014jsa ; El-Showk:2016mxr ; Esterlis:2016psv ; Li:2017agi ; Li:2017ukc ; Li:2021uki ; Kantor:2021kbx ; Kantor:2021jpz ; Afkhami-Jeddi:2021iuw ; Laio:2022ayq ; Li:2023tic . In Guo:2023qtt and the present work, we investigated the $\phi^{2n}$ and $\phi^{2n+1}$ deformations of the generalized free theory, i.e., $\phi^{m}$ interactions with integer power $m\geq 3$.  [^38] Can the generalized $\phi^{m}$ theories be realized by lattice models? A natural extension is to consider noninteger $m$. Some simple examples are the fractional power $m=\frac{m_{1}}{2},\frac{m_{1}}{3},\frac{m_{1}}{4},\dots$, where $m_{1}$ is a large enough integer. The $\phi^{m}$ deformations with irrational power or even complex power are also interesting from the perspective of analytic continuation, as they can connect the standard integer-power cases in the theory space, but we expect that they are nonunitary even for the canonical case of $k=1$. The $\mathcal{PT}$ symmetric theories are particularly relevant because they can satisfy the physical requirement of a bounded-from-below spectrum in a continuous range of power Bender:1998ke ; Bender:2007nj ; Bender:2023cem .  [^39] It is not clear to us how to capture the $\mathcal{PT}$-symmetric CFTs with even power $m$ using the $\epsilon$ expansion, as one needs to be more careful about the quantization scheme and it is not an easy task to handle multidimensional Stokes sectors, i.e., Lefschetz thimbles. See Felski:2021evi and references therein. Furthermore, it seems more natural to use the combination $i\phi$ as the building block for $\mathcal{PT}$-symmetric theories. Recently, the low-dimensional $\mathcal{PT}$ symmetric models were revisited by bootstrap methods in Li:2022prn ; Khan:2022uyz ; Li:2023nip ; John:2023him ; Li:2023ewe ; Li:2024rod , using the equations of motion in the Lagrangian or Hamiltonian formalism. It would be interesting to extend the low-dimensional bootstrap methods to higher dimensions. Along the lines of analytic continuation, one may also bootstrap the non-perturbative conformal field theory by the homotopy continuation method Li:2024ggr , where the $\epsilon$ expansion can provide regular boundary conditions near but away from the potentially dangerous Gaussian limit.  [^40] Conformal symmetry implies the existence of zero-norm states associated with specific Gaussian scaling dimensions. Usually, the zero-norm states should decouple to avoid inconsistency, such as divergences in the Gaussian limit.
@@
In addition, the $\phi^{4}$ deformation of a canonical free theory with boundaries or defects was studied using the multiplet recombination and equation of motion Yamaguchi:2016pbj ; Soderberg:2017oaa ; Dey:2020jlc ; Nishioka:2022odm ; Nishioka:2022qmj ; Bissi:2022bgu ; SoderbergRousu:2023pbe . An interesting future direction is to consider the $\phi^{m}$ deformations and higher-derivative generalizations, as well as other types of defects. Another direction would be to study CFTs on nontrivial manifolds. The case of real projective space was carried out for the $\phi^{3}$ and $\phi^{4}$ theory Hasegawa:2016piv ; Hasegawa:2018yqg . It would also be interesting to apply the multiplet-recombination method to CFTs at finite temperature.
@@ Appendix A The action of Laplacians
Here $\Delta_{i}$ denotes the scaling dimension of $\mathcal{O}_{i}$. The result of the $\Box_{x_{1}}$ action takes the form
@@
=\;= $\Bigg{[}\frac{h_{123}\left(h_{123}+h_{132}+2-d\right)}{x_{12}^{2}}+\frac{h_{132}\left(h_{123}+h_{132}+2-d\right)}{x_{13}^{2}}-\frac{h_{123}h_{132}\,x_{23}^{2}}{x_{12}^{2}x_{13}^{2}}\Bigg{]}\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}\,.$ (109)
@@
$C_{i,j}^{a,b}$ $=\bigg{(}\frac{a}{2}\bigg{)}_{i}\bigg{(}\frac{b}{2}\bigg{)}_{j}\left(\frac{a+b}{2}-\frac{d-2}{2}+i+j-k\right)_{2k-i-j}\,.$ (112)
@@
×$\hskip 128.0374pt\times B_{i_{3},j_{3}}C_{i_{3},j_{3}}^{h_{231}+2(j_{2}+k-i_{1}-j_{1}),h_{132}+2(j_{1}+k-i_{2}-j_{2})}x_{12}^{2(i_{3}+j_{3}-i_{1}-i_{2}-k)}$
×$\hskip 128.0374pt\times x_{13}^{2(i_{2}+j_{2}-j_{1}-j_{3}-k)}x_{23}^{2(i_{1}+j_{1}-j_{2}-i_{3}-k)}\Bigg{]}\frac{1}{x_{12}^{h_{123}}x_{13}^{h_{132}}x_{23}^{h_{231}}}\,.$ (113)
@@ Appendix B Analytic bootstrap at subleading twist
Let us start with the generalized Yang-Lee edge singularity. We examine the subleading term of the crossing equation (31) in the $z\rightarrow 0$ expansion. On the left-hand side of the crossing equation, both $\mathcal{J}^{(0)}_{\ell}$ and $\mathcal{J}^{(1)}_{\ell}$ contribute to the infinite sum. The lightcone expansion of the conformal blocks is written as
$G_{\mathcal{O}}(z,\bar{z})=z^{(\Delta_{\mathcal{O}}-\ell)/2}k_{h}(\bar{z})+\sum_{i=-1}^{1}c_{i}\,z^{(\Delta_{\mathcal{O}}-\ell)/2+1}k_{h+i}(\bar{z})+O(z^{(\Delta_{\mathcal{O}}-\ell)/2+2})\,.$ (114)
For $\mathcal{O}=\mathcal{J}^{(0)}_{\ell}$, the relevant part is the subleading term in this expansion. Here, $c_{i}$ are obtained by requiring that (114) solves the Casimir equation Dolan:2003hv . The results are
@@
$+z^{\Delta_{\smash{\phi}}+1}\left(\sum_{i=-1}^{1}c_{i}k_{h^{(0)}+i}(\bar{z})+\frac{1}{2}\tilde{\gamma}_{\mathcal{J}^{(0)}_{\ell}}\sum_{i=-1}^{1}c_{i}k_{h^{(0)}+i}(\bar{z})\log z+O(\tilde{\gamma}_{\smash{\mathcal{J}^{(0)}_{\ell}}}^{2})\right)+O(z^{(\Delta_{\smash{\mathcal{J}^{(0)}_{\ell}}}-\ell)/2+2})\,,$ (116)
@@
$G_{\mathcal{J}^{(1)}_{\ell}}(z,\bar{z})=z^{\Delta_{\smash{\phi}}+1}\left(k_{h^{(1)}}(\bar{z})+\frac{1}{2}\tilde{\gamma}_{\mathcal{J}^{(1)}_{\ell}}k_{h^{(1)}}(\bar{z})\log z+O\big{(}\tilde{\gamma}_{\smash{\mathcal{J}^{(1)}_{\ell}}}^{2}\big{)}\right)+O(z^{(\Delta_{\smash{\mathcal{J}^{(1)}_{\ell}}}-\ell)/2+1})\,.$ (117)
@@
The enhanced singularity on the right-hand side of the crossing equation is $\Delta_{\smash{\phi_{\text{f}}}}(\frac{1-\bar{z}}{\bar{z}})^{-\Delta_{\smash{\phi_{\text{f}}}}}$, which corresponds to the identity-operator contribution. Since $\Delta_{\smash{\phi_{\text{f}}}}=2k$, the asymptotic relation implies that the coefficient in (119) is given by $2S_{-2k}(h^{(1)})$ at order $\epsilon^{0}$. Subtracting the contribution from the leading-twist operators $\mathcal{J}^{(0)}_{\ell}$, we find the free squared OPE coefficients
@@
$P_{\phi}\frac{\Gamma(2k)}{\Gamma(k)^{2}}\Bigg{[}\left(\Delta_{\phi}+\frac{\Delta_{\phi}\left(\Delta_{\phi}-2\right)}{4}\right)\sum_{i=0}^{k-1}(-1)^{i+1}\frac{(1-2k)_{i}(k)_{i}}{i!\,(1-k)_{i}}\left(\frac{1-\bar{z}}{\bar{z}}\right)^{i-k}-\frac{\left(d-2\right)\Delta_{\phi}^{2}}{4\left(d-2\Delta_{\phi}-2\right)}$
@@
where we have only kept the terms of order $\epsilon^{1}$. Using (22), (35), and (37), we find the coefficients in the infinite sum (121). The anomalous dimensions $\tilde{\gamma}_{\smash{\mathcal{J}^{(1)}_{\ell}}}$ are obtained by subtracting the contribution from $\mathcal{J}^{(0)}_{\ell}$. The results agree with those from the multiplet recombination.
@@
which is the $m=1$ version of (78) at leading twist. For the anomalous dimensions, we consider the enhanced singularity given by (B). From this, we extract the coefficients in the infinite sum of the form (121). Here we need to account for the $N$-dependent factors in the crossing equations (D)–(D). As a result, the coefficients in the infinite sums have the following relation with those in the Yang-Lee case:
@@
Multiplying by the factor $2\sin^{2}(\pi k\,\frac{2n-3}{2n-1})$, we obtain the corresponding dDisc, as in (4.2). At subleading order in the $z\rightarrow 0$ expansion, we find the anomalous dimensions of $\mathcal{J}^{(1)}_{\ell}$ using the leading-twist anomalous dimension, as well as the free squared OPE coefficients (101) and (126). The results are consistent with those obtained from the multiplet recombination.